The cost of mailing a postcard to the USA from France is eighteen cents less than twice the cost of mailing a postcard in Germany. The total cost to mail a postcard from each country is $2.04. Find the cost to mail a postcard in each country.
Answer by checkley71(8403)

LET X BE THE PRICE FOR THE GERMAN POST CARD.
X+(2X-.18)=2.04
3X=2.04+.18
3X=2.22
X=2.22/3
X=$.74 FOR THE GERMAN POST CARD
2*74-18=1.30 FOR THE FRENCH POST CARD
.74+1.30=2.04